OpenStudy (natsent06):

If 9x = 36, what is the value of 1/4x? A) 9 B) 1 C) 144 D) 72

OpenStudy (anonymous):

b

OpenStudy (texaschic101):

9x = 36 ... we need to find x, so divide both sides by 9 x = 4 so 1/4x = 1/4(4) = 1